i really need someones help with this problem
Find the slope of the line that passes through...
(-6,-5)and(4,1)
thanks a bunch

Found 2 solutions by THANApHD, checkley71:
Answer by THANApHD(104)

you know the formula
that a slope m is equals to = (y1-y2)/(x1-x2)
where (x1,y1); (x2,y2) are two cordinates fromthe line
so the solution is m=(-5-1)/(-6-4)
= -6/-10
= 0.6

Answer by checkley71(8403)

YOU HAVE (X1,Y1) & (X2,Y2) OR (-6,-5) & (4,1) SO ALL WE NEED TO DO IS FIND
Y2-Y1 & X2-X1 SO WE HAVE FOR THE Y VALUES (1-[-5]) OR 1+5=6 FOR THE Y VALUE
FOR THE X VALUE WE HAVE (4-[-6]) OR 4+6=10
NOW WE DIVIDE THE Y VALUE BY THE X VALUE THUS 6/10 OR .6 FOR THE SLOPE OF A LINE PASSING THROUGH THE POINTS (-6,-5) & (4,1)